The area of a typical eardrum is about 5.1 × 10?5 m2

a) Find the sound power (the energy per second) incident on an eardrum at the threshold of hearing 1 × 10?12 W/m2 . Answer in units of W.

b) Find the sound power incident on an eardrum at the threshold of pain 1 W/m2 . Answer in units of
Answer #1

As we know that;

a)

Sound power incident = 10^-12 * 5.1 * 10^-5

= 5.1 * 10^-17 W

Then,
b)

Power incident = 1 * 5.1 * 10^-5

= 5.1 * 10^-5 W